Can someone please help with these?
Answer #1

1 )

using relation

T = 2\pi ( L/g )1/2

T g1/2 = constant

T1 g11/2 = T2 g21/2

6 x 9.81/2 = T2 x 1.631/2  

T2 = 14.711 sec

2 )

using conservation of energy

m g ( 2 + x ) = 1/2 k x2   

99.2 X 9.8 ( 2 + x ) = 0.5 x 1.4 x 104 x x2  

1944.32 + 972.16 x = 7000 x2

7000 x2 - 972.16 x - 1944.32 = 0

by solving

x = 0.6010 m

or

x = 60.1 cm